Given fixed cost:$40,500
Total variable cost: $192,000
Unit selling price:$100
Units sold:15,000
Find: CM
Find break-even units.
Find units required to sell if need $100,000 profit
Answer #1

Total fixed cost = $40500

Total variable cost = $192000

Units sold= 15000

Variable cost per unit =$ 192000/15000 = $12.8

Unit selling price = $100

Contribution Margin per unit (CM) = Unit sale price - Unit variable cost = $100-$12.8 = $87.2

Break even point in units = Fixed cost / CM = 40500/87.2 \approx 464

Then, Break even point in dollars = Break even point in units * Unit selling price = 464*100 = $46400

Now, number of units to be produced for desired profit = Desired profit in dollars/contribution margin per unit + Break even point in units

or, Number of units to be produced = 100000/87.2 + 464 = 1146.8 + 464 \approx 1611